Verify the following chemical equation is balanced

TiO2 + 2C → Ti + 2CO

If you begin with 15.9 kilograms of carbon, how many moles of Ti can be produced?



1
Expert's answer
2022-04-16T04:09:34-0400

m(C) = 15.9 kg = 15900 g

n(Ti) - ?

Answer:

n(Ti) = 1/2 n(C) = 1/2 m(C) / M(C)

n(Ti) = 1/2(15900g/12g/mol) = 650 mol

Answer: 650 mol